# Discovering environments with XRM

### v1

@@ Abstract
Successful out-of-distribution generalization requires environment annotations. Unfortunately, these are resource-intensive to obtain, and their relevance to model performance is limited by the expectations and perceptual biases of human annotators. Therefore, to enable robust AI systems across applications, we must develop algorithms to automatically discover environments inducing broad generalization. Current proposals, which divide examples based on their training error, suffer from one fundamental problem. These methods add hyper-parameters and early-stopping criteria that are impossible to tune without a validation set with human-annotated environments, the very information subject to discovery. In this paper, we propose Cross-Risk Minimization (XRM) to address this issue. XRM trains two twin networks, each learning from one random half of the training data, while imitating confident held-out mistakes made by its sibling. XRM provides a recipe for hyper-parameter tuning, does not require early-stopping, and can discover environments for all training and validation data. Domain generalization algorithms built on top of XRM environments achieve oracle worst-group-accuracy, solving a long-standing problem in out-of-distribution generalization.
@@ 1 Introduction
AI systems pervade our lives, spanning applications such as finance (Hand and Henley, 1997), healthcare (Jiang et al., 2017), self-driving vehicles (Bojarski et al., 2016), and justice (Angwin et al., 2016). While machines appear to outperform humans on such tasks, these systems fall apart when deployed in testing conditions different to their experienced training environments (Geirhos et al., 2020). For instance, during the COVID-19 pandemic, it was shown that thoracic x-ray classifiers latched onto spurious correlations—such as patient age, scanning position, or text fonts—as shortcuts to minimize their own training error (Heaven, 2021). This resulted in “an alarming situation in which the systems appear accurate, but fail when tested in new hospitals” (DeGrave et al., 2021).
@@
Generally speaking, AI systems perform worse on groups of examples under-represented in the training data (Barocas et al., 2019). To drive this point home, consider the left side of figure 1, illustrating the Waterbirds problem (Sagawa et al., 2019). This task considers two class labels, landbirds and waterbirds, collected in two landscape environments, land and water. These combine into four groups: a majority group of waterbirds in water (73% of training examples), landbirds in land (22%), waterbirds in land (4%), and a minority group of landbirds in water (1%). On this problem, learning machines latch onto the landscape spurious feature, because it can separate the majority of examples, exhibits a larger signal-to-noise ratio, and yields the maximum-margin classifier amongst those with zero training error. For instance, consider an empirical risk minimization (Vapnik, 1998, ERM) baseline, which ignores environment information. As shown in the right panel of figure 1, ERM results in a worst-group-accuracy of $61\%{}$—attained, in fact, on the minority group.
@@
To improve upon ERM, researchers have developed a myriad of domain generalization (DG) algorithms (Zhou et al., 2022; Wang et al., 2021). These methods consider environment annotations to uncover invariant (environment-generic) patterns and discard spurious (environment-specific) correlations (Arjovsky et al., 2019). As figure 1 shows, the DG algorithm group distributionally robust optimization (Sagawa et al., 2019, GroupDRO) achieves a worst-group-accuracy of $87\%{}$. This outperforms ERM by over twenty five points, a sizeable gap!
@@
Figure 1: (a) Waterbirds classification problem, containing four groups: a majority group of waterbirds in water, landbirds in land, waterbirds in land, and a minority group of landbirds in water. Learning machines latch onto spurious landscape features, as it allows to classify the majority of examples, exhibits a larger signal-to-noise ratio, and yields the maximum-margin classifier amongst solutions with zero training error. (b) Worst-group-accuracy (minority-group-accuracy) for different methods. (Dotted line) An ERM baseline ignoring group annotations achieves $61\%{}$. (Dashed line) The GroupDRO domain generalization method with human group annotations, considered our oracle, achieves $87\%{}$. (Dashdot blue line) Prior work to discover groups requires early-stopping with surgical precision. This is an impossible ask without a validation set with human group annotations, the very information subject to discovery. Realistically, these methods would converge at $68\%{}$. (Solid red line) Our proposed XRM enables an oracle performance of $87\%{}$ at convergence.
@@
While promising, DG algorithms require environment annotations. These are resource-intensive to obtain, and their relevance to downstream model performance is limited by the expectations, precision, and perceptual biases of human annotators. Moreover, no fixed environment annotations that we may choose and revise can serve the idiosyncrasies of all the different DG algorithms. In extremis, and by virtue of the combinatorial-explosion of ways in which two examples can be similar or dissimilar (Goodman, 1972), the patterns that deceive a learning system could be alien or invisible to our human eyes (Goodfellow et al., 2014). Because of these reasons, robust AI systems are currently confined to small data collections, and their promise in the large-scale setting remains unfulfilled.
@@
In light of the above, some researchers began developing algorithms for the automatic discovery of environments from data (Bao and Barzilay, 2022; Zheran Liu et al., 2021; Lahoti et al., 2020; Dagaev et al., 2021; Creager et al., 2020; Nam et al., 2020). In broad strokes, these methods build a robust system in two phases. In phase-1, these methods learn a label predictor to distribute training examples in two environments, based on their training error. In phase-2, a DG algorithm is trained on top of the discovered environments to produce the robust system of interest. Unfortunately, this pipeline suffers from one fundamental issue. In particular, phase-1 needs to control the capacity of the label predictor with surgical precision, such that the discovered environments differ only in spurious correlation. As illustrated in figure 1, these methods discover environments that yield strong phase-2 systems only at the knife’s edge of very precisely early-stopped phase-1 training iterations—yielding a decent but still worse than oracle performance of $79\%{}$—and otherwise converge to a worst-group-accuracy of $68\%{}$. Unfortunately, we lack a signal for such a fine call, so methods for environment discovery resort to a validation set with human environment annotations. In practice, this wraps phase-1 and phase-2 with a cross-validation envelope to directly select a model maximizing validation worst-group-accuracy. Alas, at least in our view, this defeats the raison d’être of environment discovery.
@@
We propose Cross-Risk Minimization (XRM), a simple method for environment discovery that requires no human environment annotations whatsoever. XRM trains two twin label predictors, each holding-in one random half of the training data. During training, XRM instructs each twin to imitate confident held-out mistakes made by their sibling. This results in an “echo-chamber” where twins increasingly rely on bias, converging on a pair of environments that differ in spurious correlation, and share the invariances that fuel downstream out-of-distribution generalization. After twin training, a simple cross-mistake formula allows XRM to annotate all of the training and validation examples with environments. As our experiments show, XRM endows DG algorithms with oracle-like performance across benchmarks, solving a long-standing problem in out-of-distribution generalization. Returning one final time to figure 1, we observe that XRM+GroupDRO converges to $87\%{}$ worst-group-accuracy on Waterbirds, matching the oracle!
The sequel is as follows. Section 2 reviews the setup of DG, pointing-out the need for environment annotations as one of its major shortcomings. Section 3 surveys prior works on environment discovery, their issues surrounding hyper-parameter tuning and early-stopping, and their consequent reliance on validation sets with human environment annotations. Section 4 introduces XRM, a novel objective to discover machine-tailored environments for all training and validation data. Experiments in Section 5 show that DG algorithms built on top of XRM environments achieve oracle-like performance, and Section 6 closes with thoughts for future work.
2 Learning invariances across known environments
In domain generalization (DG), our goal is to build learning systems that perform well beyond the distribution of the training data. To this end, we collect examples under multiple training environments. Then, DG algorithms search for patterns that are invariant across these training environments—more likely to hold during test time—while discarding environment-specific spurious correlations (Arjovsky et al., 2019). More formally, we would like to learn a predictor $f$ to classify inputs $x$ into their appropriate labels $y$, and across all relevant environments $e\in\mathcal{E}$:
@@ 2 Learning invariances across known environments
In practical applications, the DG problem (1) is under-specified in two important ways. Firstly, we only get to train on a subset of all of the relevant environments $\mathcal{E}$, called the training environments $\mathcal{E}_{\text{tr}}\subset\mathcal{E}$. Yet, the quality of our predictor continues to be the worst classification accuracy across all environments $\mathcal{E}$. Secondly, and for each training environment $e\in\mathcal{E}_{\text{tr}}$, we do not observe its entire data distribution $P^{e}$, but only a finite set of iid examples $(x_{i},y_{i},e_{i}=e)$. In sum, the information at our disposal to address the DG problem (1) is the training dataset $\mathcal{D}_{\text{tr}}=\{(x_{i},y_{i},e_{i})\}_{i=1}^{n}$, where $(x_{i},y_{i})$ is an input-label pair drawn from the distribution $P^{e_{i}}$ associated to the training environment $e_{i}\in\mathcal{E}_{\text{tr}}$. Armed with $\mathcal{D}_{\text{tr}}$, we approximate (1) by the observable optimization problem
$f\in\mathop{\mathrm{argmin}}_{\tilde{f}}\sup_{e\in\mathcal{E}_{\text{tr}}}R_{n}^{e}(\tilde{f}),$ (2)
where $R^{e}_{n}(f)=\frac{1}{|\mathcal{D}^{e}_{\text{tr}}|}\sum_{(x_{i},y_{i})\in\mathcal{D}^{e}_{\text{tr}}}\ell(f(x_{i}),y_{i})$ is the empirical risk (Vapnik, 1998) across the data $\mathcal{D}^{e}_{\text{tr}}=\{(x_{i},y_{i})\in\mathcal{D}_{\text{tr}}:e_{i}=e\}$ from the training environment $e\in\mathcal{E}_{\text{tr}}$.
Environments and groups
In its full generality, domain generalization is an admittedly daunting task. To alleviate the burden, much prior literature considers the simplified version of group shift (Sagawa et al., 2019). The problem formulation is equivalent: we observe each input $x_{i}$ together with some attribute $e_{i}$ and label $y_{i}$, and define one group $g\equiv e\times y$ per attribute-label combination. (We use the terms “environment” and “attribute” interchangeably.) Again, we assume access to one training set of triplets $(x_{i},y_{i},e_{i})$ to learn our predictor, and one similarly formatted validation set available for hyper-parameter tuning and model selection purposes. Next, we put in place one important simplifying assumption $\mathcal{E}=\mathcal{E}_{\text{tr}}$, namely no new environments appear during test time. Consequently, the quality of our predictor can be directly estimated as the worst-group-accuracy in the validation set. Because most learning algorithms focus on minimizing average training error, oftentimes the worst-group-accuracy happens to be on the group with the smallest number of examples, also known as the minority group.
In practice, different DG algorithms (Gulrajani and Lopez-Paz, 2020; Zhou et al., 2022; Wang et al., 2021; Yang et al., 2023) target different types of invariance, learned across training environments $\mathcal{E}_{\text{tr}}$, assumed to hold across testing environments $\mathcal{E}_{\text{te}}$, and implemented as various innovations to the objective (2). As discussed in section 1, some DG algorithms outperform by a large margin methods ignoring environment (cf. attribute, group) information, such as ERM.
Despite their promise, the main roadblock towards large-scale domain generalization is their reliance on humanly annotated environments, attributes, or groups. These annotations are resource-intensive to obtain. Moreover, the expectations, precision, and perceptual biases of annotators can lead to environments conducive of sub-optimal out-of-distribution generalization. Different machine learning models fall prey to different kinds of spurious correlations. In addition, there are plenty of complex and subtle interactions between environment definition, function class, distributional shift, and cultural viewpoint (Lopez-Paz et al., 2022). Therefore, environment annotations are helpful only when revealing spurious and invariant patterns under the lens of the learning system under consideration. Could it be possible to design algorithms for the automatic discovery of environments tailored to the learning machine and data at hand?
3 Discovering environments
@@
Let us reconsider the problem of domain generalization without access to environment annotations. This time it suffices to talk about one training distribution $P^{\text{tr}}$ and one testing distribution $P^{\text{te}}$. Our training data is a collection of input-label pairs $(x_{i},y_{i})$, each drawn iid from the training distribution. While the training distribution $P^{\text{tr}}$ may be the mixture of multiple environments describing interesting invariant and spurious correlations, this rich heterogeneity is shuffled together and unbeknown to us. But, if we could “unshuffle” the training distribution and recover the environments therein, we could invoke the domain generalization machinery from the previous section and hope for a robust predictor. This is the purpose of automatic environment discovery.
To discover environments and learn from them, most prior work implements a pipeline with two phases. On phase-1, train a label predictor and distribute each training example into two environments, depending on whether the example is correctly or incorrectly classified. On phase-2, train a DG algorithm on top of the discovered environments. Crucially, one must control the capacity of the label predictor in phase-1 with surgical precision, such that it relies only on prominent, easier-to-learn spurious correlations. If the environments discovered in phase-1 differ only in spurious correlation, as we would like, then the DG algorithm from phase-2 should be able to zero-in on invariant patterns more likely to generalize to the test distribution $P^{\text{te}}$. On the unlucky side, if phase-1 produces a zero-training-error predictor, we would be providing phase-2 with one non-vacuous, non-informative environment—the training data itself!
@@
As a result, proposals for environment discovery differ mainly in how to control the capacity of the phase-1 label predictor. For example, the too-good-to-be-true prior (Dagaev et al., 2021) employs a predictor with a small parameter count. Just train twice (Zheran Liu et al., 2021, JTT) and environment inference for invariant learning (Creager et al., 2020, EIIL) train a phase-1 predictor for a limited number of epochs. Learning from failure (Nam et al., 2020, LfF) biases the predictor towards the use of “simple” features by applying a generalized version of the cross entropy loss. Other proposals, such as learning to split (Bao and Barzilay, 2022, LS) and adversarial re-weighted learning (Lahoti et al., 2020, ARL) complement capacity control with adversarial games.
However regularized, all of these methods suffer from one fundamental problem. More specifically, these phase-1 strategies add hyper-parameters and early-stopping criteria, but remain silent on how to tune them. As illustrated in figure 1 for Waterbirds, methods like the above discover environments leading to competitive generalization only when phase-1 is trained for a number of iterations that fall within a knife’s edge. Quickly after that, the performance of the resulting phase-2 system falls off a cliff, landing at ERM-like worst-group-accuracy.
Prior works keep away from this predicament by assuming a validation set with human environment annotations. Then, it becomes possible to simply wrap phase-1 and phase-2 into a cross-validation pipeline that promotes validation worst-group-accuracy. Alas, this defeats the entire purpose of environment discovery. In fact, if we have access to a small dataset with human environment annotations, these examples suffice to fine-tune the last layer of a deep neural network towards state-of-the-art worst-group-accuracy (Izmailov et al., 2022). Looking forward, could we develop an algorithm for environment discovery that requires no human annotations whatsoever, and robustly yields oracle-like phase-2 performance?
@@ 4 Cross-Risk Minimization (XRM)
We propose Cross-Risk Minimization (XRM), an algorithm to discover environments without the need of human supervision. XRM comes with batteries included, namely a recipe for hyper-parameter tuning and a formula to annotate all training and validation data. As we will show in section 5, environments discovered by XRM endow phase-2 DG algorithms with oracle performance.
The blueprint for phase-1 with XRM is as follows. XRM trains two twin label predictors, each holding-in one random half of the training data (section 4.1). During training, XRM biases each twin to absorb spurious correlation by imitating confident held-out mistakes from their sibling (section 4.2). XRM chooses hyper-parameters for the twins based on the number of imitated mistakes (section 4.3). Finally, and given the selected twins, XRM employs a simple “cross-mistake” formula to discover environment annotations for all of the training and validation examples (section 4.4). Algorithm 1 serves as a companion to the descriptions below; appendix D contains a real PyTorch implementation. The runtime of phase-1 with XRM is akin to one ERM baseline on the training data.
@@
Output: discovered environments for training $\{e_{i}\}_{i=1}^{n}$ and validation $\{\tilde{e}_{i}\}_{i=1}^{m}$ examples
@@
  • Init two label predictors $f^{a}$ and $f^{b}$ at random; calibrate softmax temperatures on held-in data
@@
    • Compute held-in softmax predictions $p^{\text{in}}_{i}=m^{a}_{i}f^{a}(x_{i})+m_{i}^{b}f^{b}(x_{i})$
    • Compute held-out softmax predictions $p^{\text{out}}_{i}=m^{b}_{i}f^{a}(x_{i})+m^{a}_{i}f^{b}(x_{i})$
    • Update $f^{a}$ and $f^{b}$ to minimize the class-balanced held-in cross-entropy loss $\ell(p^{\text{in}},y)$
    • Flip $y_{i}$ into $y^{\text{out}}_{i}=\mathrm{argmax}_{j}p^{\text{out}}_{i,j}$, with prob. $(p^{\text{out}}_{i,y^{\text{out}}_{i}}-1/n_{\text{classes}})\cdot n_{\text{classes}}/(n_{\text{classes}}-1)$

### v2

@@ Abstract
Environment annotations are essential for the success of many out-of-distribution (OOD) generalization methods. Unfortunately, these are costly to obtain and often limited by human annotators’ biases. To achieve robust generalization, it is essential to develop algorithms for automatic environment discovery within datasets. Current proposals, which divide examples based on their training error, suffer from one fundamental problem. These methods introduce hyper-parameters and early-stopping criteria, which require a validation set with human-annotated environments, the very information subject to discovery. In this paper, we propose Cross-Risk Minimization (XRM) to address this issue. XRM trains twin networks, each learning from one random half of the training data, while imitating confident held-out mistakes made by its sibling. XRM provides a recipe for hyper-parameter tuning, does not require early-stopping, and can discover environments for all training and validation data. Algorithms built on top of XRM environments achieve oracle worst-group-accuracy, addressing a long-standing challenge in OOD generalization. Code available at https://github.com/facebookresearch/XRM.
Machine Learning, ICML
@@ 1 Introduction
AI systems pervade our lives, spanning applications such as finance (Hand & Henley, 1997), healthcare (Jiang et al., 2017), self-driving vehicles (Bojarski et al., 2016), and justice (Angwin et al., 2016). Despite outperforming humans in many tasks, these systems fall apart under testing conditions different from their training environments (Geirhos et al., 2020). For instance, during the COVID-19 pandemic, thoracic x-ray classifiers incorrectly latched onto spurious correlations such as patient’s age or position (Heaven, 2021), leading to “an alarming situation in which the systems appear accurate, but fail when tested in new hospitals” (DeGrave et al., 2021).
@@
Generally speaking, AI systems underperform on under-represented groups in training data (Barocas et al., 2019). The Waterbirds problem (Sagawa et al., 2019), depicted in Figure 1, illustrates this with two classes (landbirds and waterbirds) in two landscape environments (land and water), forming four groups: a majority group of waterbirds in water (73% of examples), landbirds in land (22%), waterbirds in land (4%), and a minority group of landbirds in water (1%). On this problem, learning machines often favor the landscape spurious feature to classify the majority of examples and memorizes the remaining minorities to achieve zero training error. An empirical risk minimization (ERM) baseline, ignoring environment data (Vapnik, 1998), achieves a mere 61% worst-group-accuracy, specifically on the minority group, as illustrated in Figure 1’s right panel.
@@
To improve upon ERM, researchers have developed a myriad of OOD generalization algorithms (Zhou et al., 2022a; Wang et al., 2021). These methods use environment annotations to uncover invariant (environment-generic) patterns and discard spurious (environment-specific) ones (Arjovsky et al., 2019). As Figure 1 shows, group distributionally robust optimization (Sagawa et al., 2019, GroupDRO) achieves a worst-group-accuracy of $87\%{}$. This outperforms ERM by over twenty five points, a sizeable gap!
@@
Figure 1: (a) Waterbirds problem with four groups: a majority group of waterbirds in water, landbirds in land, waterbirds in land, and a minority group of landbirds in water. Models often rely on spurious features to classify the majority of examples and then memorize the minority examples. (b) Worst-group-accuracy (minority) for different methods. (Dotted line) ERM achieves $61\%{}$. (Dashed line) GroupDRO with human group annotations (oracle) achieves $87\%{}$. (Dashdot blue line) Prior work to discover groups requires early-stopping with surgical precision. (Solid red line) XRM enables an oracle performance of $87\%{}$ without requiring early stopping.
@@
While promising, OOD algorithms targeting sub-population shift require environment annotations, which are costly to obtain and limited by human annotators’ biases and precision. Moreover, no single set of environment annotations fits all OOD algorithms. The patterns misleading a learning system might be alien or invisible to humans (Goodfellow et al., 2014). Because of these reasons, OOD generalization is currently confined to small data collections, and their promise in the large-scale setting remains unfulfilled.
@@
In light of the above, researchers developed algorithms for automatic environment discovery (Bao & Barzilay, 2022; Zheran Liu et al., 2021; Zhang et al., 2022b; Lahoti et al., 2020; Dagaev et al., 2021; Creager et al., 2020; Nam et al., 2020). These methods typically build a robust system in two phases. In phase-1, these methods train a classifier to categorize training examples in two environments, based on their training error. In phase-2, an OOD algorithm is trained on top of the discovered environments. However, this pipeline suffers from one fundamental issue: the need for precisely controlling the classifier’s capacity, ensuring that the discovered environments differ only in spurious correlations. As Figure 1 shows, peak performance (79%) is reached with exact early-stopping in phase-1; without it, accuracy drops to 68%. Lacking a precise indicator of early-stopping, environment discovery methods depend on validation sets with human annotations. Alas, at least in our view, this defeats the raison d’être of environment discovery.
@@
We propose Cross-Risk Minimization (XRM), a simple method for environment discovery that requires no human environment annotations whatsoever. XRM trains two twin networks, each holding-in one random half of the training data. During training, XRM instructs each twin to imitate confident held-out mistakes made by their sibling. This results in an “echo-chamber” where twins increasingly rely on bias, converging on a pair of environments that differ in spurious correlation, and share the invariances that fuel downstream out-of-distribution generalization. After twin training, a simple cross-mistake formula allows XRM to annotate all of the training and validation examples with environments. As our experiments show, XRM endows OOD generalization algorithms with oracle-like performance across benchmarks. Returning one final time to Figure 1, we observe that XRM+GroupDRO converges to $87\%{}$ worst-group-accuracy on Waterbirds, matching the oracle!
The sequel is as follows. Section 2 details the problem formulation. Section 3 surveys prior research on environment discovery. Section 4 describes XRM. Section 5 demonstrates the effectiveness of XRM, while Section 6 discusses its potential shortcomings. Section 7 concludes with thoughts for future work.
2 Learning Invariances Across Environments
The goal of OOD generalization is to build learning systems that perform well beyond the training data distribution. To this end, we collect examples from multiple environments and OOD algorithms search for invariant patterns across these environments, while disregarding environment-specific spurious correlations (Arjovsky et al., 2019). Formally, we seek a predictor $f$ that classifies inputs $x$ into labels $y$, across all relevant environments $e\in\mathcal{E}$:
@@ 2 Learning Invariances Across Environments
In its full generality, OOD generalization in (1) is an admittedly daunting task. To alleviate the burden, prior literature often considers the simplified and more practical version of sub-population shift (Sagawa et al., 2019). Given a dataset $\mathcal{D}=\{(x_{i},y_{i},e_{i})\}_{i=1}^{n}$, the supremum in (1) is replaced by a maximum over the training environments and the risk for each environment is approximated by the empirical risk (Vapnik, 1998). The effectiveness of an OOD algorithm is then assessed by its worst-group-accuracy on a validation set.
In practice, several OOD algorithms have been successful in learning invariances across environments  (Gulrajani & Lopez-Paz, 2020; Zhou et al., 2022a; Wang et al., 2021; Yang et al., 2023). Despite their promise, their large-scale application is hindered by the need for human-annotated environments, which are resource-intensive and might be even sub-optimal. Different machine learning models fall prey to different kinds of spurious correlations. In addition, there exists complex interactions between environment definition, function class, distributional shift, and cultural viewpoints (Lopez-Paz et al., 2022). Therefore, environment annotations are helpful only when revealing spurious and invariant patterns under the lens of the learning system under consideration. Could it be possible to design algorithms for the automatic discovery of environments tailored to the learning machine and data at hand?
2.1 Discovering Environments
@@
Let us reconsider the problem of OOD generalization without access to environment annotations. This time, it suffices to talk about one training distribution $P^{\text{tr}}$ and one testing distribution $P^{\text{te}}$. Our training data is a collection of input-label pairs $(x_{i},y_{i})$, each drawn iid from the training distribution. While $P^{\text{tr}}$ may be the mixture of multiple environments describing interesting invariant and spurious correlations, this rich heterogeneity is shuffled together and unbeknown to us. But, if we could “unshuffle” the training distribution and recover the environments therein, we could invoke the OOD generalization machinery from the previous section and hope for a robust predictor. This is the purpose of automatic environment discovery.
3 Related Work on Environment Discovery
To discover environments, prior work often train a classifier and then assign each training example to two environments based on their loss or classification accuracy. Crucially, one must control the capacity of the classifier with surgical precision, such that it relies only on the spurious correlations. It is only in such cases that the subsequent OOD generalization algorithms can successfully disregard these spurious features.
@@ 3 Related Work on Environment Discovery
As a result, proposals for environment discovery differ mainly in how to control the capacity of the classifier. For example, the too-good-to-be-true prior (Dagaev et al., 2021) employs a classifier with a small parameter count while correct-n-contrast (Zhang et al., 2022b, CnC) applies strong weight decay regularization. Just train twice (Zheran Liu et al., 2021, JTT) and environment inference for invariant learning (Creager et al., 2020, EIIL) train a classifier for a limited number of epochs. Learning from failure (Nam et al., 2020, LfF) biases the classifier towards the use of “simple” features by applying a generalized version of the cross entropy loss. Other proposals, such as learning to split (Bao & Barzilay, 2022, LS) and adversarial re-weighted learning (Lahoti et al., 2020, ARL) complement capacity control with adversarial games.
However, all these methods assume having access to a human-annotated validation set to conduct such precise capacity controls. This defeats the purpose of environment discovery. In fact, if we have access to a small dataset with human-annotated environments, these examples suffice to fine-tune the last layer of a deep network towards state-of-the-art worst-group-accuracy (Izmailov et al., 2022).
For a more detailed discussion and related work, please refer to Appendix A.
@@ 4 Cross-Risk Minimization (XRM)
We propose Cross-Risk Minimization (XRM), an algorithm to discover environments without the need of human supervision. XRM comes with batteries included, namely a recipe for hyper-parameter tuning and a formula to annotate all training and validation data. As we will show in Section 5, environments discovered by XRM endow OOD generalization algorithms with oracle performance.
The blueprint for XRM is as follows. XRM trains two twin classifiers, each holding-in one random half of the training data (Section 4.1). During training, XRM biases each twin to absorb spurious correlation by imitating confident held-out mistakes from their sibling (Section 4.2). XRM chooses hyper-parameters for the twins based on the number of imitated mistakes (Section 4.3). Finally, and given the selected twins, XRM employs a simple “cross-mistake” formula to discover environment annotations for all of the training and validation examples (Section 4.4). Algorithm 1 serves as a companion to the descriptions below; Appendix C contains a PyTorch implementation. The runtime of XRM is akin to one ERM baseline on the training data.
@@
Output: discovered environments $\{e_{i}\}_{i=1}^{n}$ and $\{\tilde{e}_{i}\}_{i=1}^{m}$
@@
  • Initialize two classifiers $f^{a}$ and $f^{b}$
@@
    • Compute held-in softmax predictions:
      $p^{\text{in}}_{i}=m^{a}_{i}f^{a}(x_{i})+m_{i}^{b}f^{b}(x_{i})$
    • Compute held-out softmax predictions:
      $p^{\text{out}}_{i}=m^{b}_{i}f^{a}(x_{i})+m^{a}_{i}f^{b}(x_{i})$
    • Update $f^{a}$ and $f^{b}$ to minimize the held-in class-balanced cross-entropy loss $\ell(p^{\text{in}},y)$
    • Flip $y_{i}$ into $y^{\text{out}}_{i}=\mathrm{argmax}_{j}p^{\text{out}}_{i,j}$, with probability:
      $(p^{\text{out}}_{i,y^{\text{out}}_{i}}-1/n_{\text{classes}})\cdot n_{\text{classes}}/(n_{\text{classes}}-1)$